OpenStudy (chris215):

Find f '(x) for f(x) = 2x3 - 4x2 + 6x - 25

OpenStudy (jojokiw3):

Do you know what to do first?

OpenStudy (chris215):

no

OpenStudy (whpalmer4):

Plug it into that same formula I gave you: \[f'(x) = \lim_{h\rightarrow 0} \frac{f(x+h)-f(x)}{h}\]and grind through the algebra

OpenStudy (jojokiw3):

Okay so first, I will give you the equation for figuring out derivatives: \[f'(x) = \lim_{h \rightarrow 0}\frac{ f(x+h)- f(x) }{ h }\]

OpenStudy (jojokiw3):

Get it right this way and I'll teach you the shortcut method. :)

OpenStudy (whpalmer4):

as an example, I'll do \(f(x) = 2x+3\) \[f'(x) = \lim_{h\rightarrow 0}\frac{(2(x+h)+3) - (2(x)+3)}{h}\]\[=\lim_{h\rightarrow 0}\frac{(2x+2h+3) - (2x+3)}{h} = \lim_{h\rightarrow 0}\frac{2x+2h+3 - 2x-3}{h} \]\[=\lim_{h\rightarrow 0}\frac{2h}{h} = 2\]

OpenStudy (chris215):

ok thanks i got f '(x) = 6x^2 - 8x - 19

OpenStudy (mathmale):

Where'd that 19 come from? Remember that the derivative of a constant, such as of -25, is zero (0).

OpenStudy (chris215):

oh opps f '(x) = 6x^2 - 8x + 6

OpenStudy (mathmale):

A lot better! :)

OpenStudy (anonymous):

perfect :)

OpenStudy (anonymous):

remember : \(\rm If~y=x^n~then~y'=n \cdot~x^{n-1}\)

OpenStudy (mathmale):

The method used here involves several rules of differentiation: 1. The constant coefficient rule 2. The power rule (given by secret88, above) 3. The addition and subtraction rules for differentiation Note that it's also possible to obtain the derivative using the definition of the derivative.

OpenStudy (chris215):

thanks :)

OpenStudy (whpalmer4):

I assumed (perhaps incorrectly) that as your previous question involved the definition of the derivative, this one was expected to be done that way instead of by the power rule. Power rule certainly involves less algebra to do!